OpenStudy (anonymous):

on a rectangler what is the area? 2x^2 by 7x^2 is it 14^2

OpenStudy (anonymous):

can someone help me

OpenStudy (anonymous):

help.....

OpenStudy (lgbasallote):

you multiply them \[\huge 2x^2 \times 7x^2 \implies (2\times 7)(x^2 \times x^2)\] \[\huge \implies (2\times 7)(x^{2+2})\] does that help?

OpenStudy (anonymous):

is that 14

OpenStudy (lgbasallote):

no

OpenStudy (lgbasallote):

hat is \[\huge x^{2+2}\]

OpenStudy (anonymous):

that is the answer

OpenStudy (anonymous):

hello

OpenStudy (lgbasallote):

NO...